if a person sell his car in 50000 rupees and got 17% profit.what will be the profit in percentage if he he sold this car in 47000 rupees
.
kindly solve this?

Answer by dkppathak(439)

if a person sell his car in 50000 rupees and got 17% profit.what will be the profit in percentage if he he sold this car in 47000 rupees
solution
profit =17 %
sp of car = Rs 50000
cp =sp/1+.profit percentage
cp 50000/1+.17
cp= 50000/1.17= Rs 42735.04
now he wants to sell at Rs 47000
new sp= Rs 47000
profit sp-cp =47000-42735.04= Rs 4264.96
profit %= profitx100/cp
profit%=4264.96x100/42735.04=9.98 %
new profit %=9.98